A circular grassy region with a 96 ft diameter has a 12 ft x 17 ft rectangular pond and its center if a 2ft concrete border is around the pond how many Square feet of concrete will be laid
Answer by ikleyn(52779)

.
The larger rectangle has dimensions 12 + 2*2 = 16 ft (width) and 17+2*2 = 21 ft (length).


The area of the concrete border is 16*21 - 12*17 = 132 square feet.

Solved.

Notice that the info about "A circular grassy region with a 96 ft diameter" is excessive, unnecessary and is not used in the solution of this problem.